kkumrimaedo:

What will be the coordinates of parallelogram ABCD after a dilation about the origin by a scale factor of 0.5? A. A'(2, 2), B'(2, 0), C'(−1, −1), D'(1, −1) B. A'(−2, 2), B'(0, 2), C'(1, −1), D'(−1, −1) C. A'(−4.5, 4.5), B'(0.5, 4.5), C'(2.5, −2.5), D'(−2.5, −2.5) D. A'(−8, 8), B'(0, 8), C'(4, −4), D'(−4, −4)

kkumrimaedo:

1 attachment
kkumrimaedo:

@Razor or @Luminexgirl just someone help T^T

kkumrimaedo:

i am very bad at all this soo i don't understand anything!

Razor:

Looking at it right now, sorry if I just look like I'm lurking. But the formula you should use for this problem is SF = X2/Y2 / X1/Y1 (I'm pretty sure. y = Cx also works, where C is the scale factor for x. A = (-4, 4) - B = (0, 4) - C = (2, -2) - and D = (-2, -2) As long as we know that the two shapes are similar we just divide the width of one by our scale factor - which is 0.5

kkumrimaedo:

so its either a or d ... *-*

Razor:

6 is the width 4 is the length Divide those both by 0.5 and then take that and slap it on a grid

Razor:

literally just take your fist and punch it on a grid

kkumrimaedo:

it just gave me 1.3333333333 -w-

kkumrimaedo:

im sounding very dumb T^T

Razor:

No you don't that was my b. I'm not very good at math but I'm trying. ¯\_(ツ)_/¯

kkumrimaedo:

its fine -w-